OpenStudy (honeymonster):

i have struggled with this for a while so please help..how to make x the subject of the formula posted in replies..

OpenStudy (anonymous):

http://openstudy.com/study#/updates/503cf88be4b04bebe0425b57 please help

OpenStudy (honeymonster):

\[\frac{ 1 }{ x+1 } - \frac{ 1 }{ y-1 } =1\]

OpenStudy (anonymous):

\[1\div \left( x+1 \right)=1\div \left( y-1 \right)+1\]

OpenStudy (phi):

can you add the right side in \[ \frac{1}{x+1}= 1+\frac{1}{y-1} \] ?

OpenStudy (phi):

by using a common denominator of y-1

OpenStudy (honeymonster):

possibly..not completely sure..sorry

OpenStudy (phi):

try. multiply the 1 by (y-1)/(y-1) then add the tops and keep the same bottom

OpenStudy (anonymous):

\[1=(x+1)[\frac{ 1 }{ y-1 }+1]\] \[\frac{ 1 }{ \frac{ 1 }{ y-1 }+1 }=x+1\] \[\frac{ 1 }{ \frac{ 1 }{ y-1 }+1 }-1=x\]

OpenStudy (anonymous):

I'm pretty sure that's correct.

OpenStudy (anonymous):

And you couldn't use y-1 as a common denominator unless you made all the elements of the equation have a denominator of y-1

OpenStudy (honeymonster):

ok thank you everybody :)